OpenStudy (anonymous):

A grid shows the positions of a subway stop and your house. The subway stop is located at (-5, 2), and your house is located at (-9, 9). What is the distance, to the nearest unit between your house and the subway stop? A. 5 B. 13 C. 8 D. 18

OpenStudy (anonymous):

helps to draw it out if you haven't

OpenStudy (anonymous):

you want the distance in between

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -5}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ -9}}\quad ,&{\color{blue}{ 9}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

 Realize that the "distance formula" is the Pythagorean Theorem: a^2 + b^2 = c^2 can you find the lengths of a and b? (subtract the x and y coordinates of both respectively)

OpenStudy (anonymous):

@Euler271 -3, 0?

OpenStudy (anonymous):

the length of a is |-9 - (-5)| and the length of b is |9 - 2|

OpenStudy (anonymous):

c = sqrt(4^2 + 7^2)

OpenStudy (anonymous):

@Euler271 so it would be 14, 7?

OpenStudy (anonymous):

they want the distance (to the nearest unit. means you have to round up or down depending on the number) its sqrt(16+49) = sqrt(65). idk what that equals without a calculator (8. something)

OpenStudy (anonymous):

@Euler271 so it would round to 8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much! you're a life saver (